each participant in a community garden is allotted a rectangular plot that meausres 18 ft. by 45 ft. how much is needed to enclose wach plot.

Answer by stanbon(75887)

each participant in a community garden is allotted a rectangular plot that meausres 18 ft. by 45 ft. how much is needed to enclose wach plot.
-----
Perimeter = 2(L + W)
P = 2(18 + 45)
---
P = 2(63)
---
P = 126 feet
=======================
Cheers,
Stan H.